OpenStudy (anonymous):

sin(theta)=1/5 whats sec (theta) and tan (theta)?

OpenStudy (jdoe0001):

is the angle in any specific quadrant?

OpenStudy (luigi0210):

Make a triangle.. \[\LARGE sin\theta=\frac{opposite}{hyp}\]

OpenStudy (luigi0210):

So..  Use \(\large \color{green}{ a^2+b^2=c^2}\) to solve for the missing side.

OpenStudy (anonymous):

thanks and no specific